An investor is in the 32 percent tax bracket and pays long-term capital gains taxes of 15 percent. What are the taxes owed (or saved in the case of losses) in the current tax year for each of the following situations?

a) Net short-term capital gains of $3,000; net long-term capital gains of $4,000

b) Net short-term capital gains of $3,000; net long-term capital losses of $4,000

c) Net short-term capital losses of $3,000; net long-term capital gains of $4,000

d) Net short-term capital gains of $3,000; net long-term capital losses of $2,000

e) Net short-term capital losses of $4,000; net long-term capital gains of $3,000

f) Net short-term capital losses of $1,000; net long-term capital losses of $1,500

g) Net short-term capital losses of $3,000; net long-term capital losses of $2,000
Answer #1

capital loss is & uso. solutim (as short term capital gain tax payable 3000 x 32% 960 longterm capital gain tax payable 4000Xd) Net capital gain = 3800+(-2000) = 1000 capital gain fax payable = 1000x 32) 320 2000 x 32%. Tax saved 2640 The Prvester haThe total tax saved is s 960 in the lissent year. 3 Net capital dess = -1000 +(-1900) -2500 Total tone saved = 2500x32% =$800